OpenStudy (anonymous):

Is the sum from 1 to infinity of n/(n+2) convergent or divergent?

OpenStudy (anonymous):

The answer is convergent, but why?

OpenStudy (amistre64):

does the limit of the sequence goes to zero ?

OpenStudy (anonymous):

No, it goes to 1. Yet the answer on the test is convergent.

OpenStudy (amistre64):

its not convergent

OpenStudy (amistre64):

\[\frac{n}{n+2}\] \[\frac{n+2-2}{n+2}\] \[\frac{n+2}{n+2}-\frac{2}{n+2}\] \[1-\cancel{\frac{2}{n+2}}^{~0}=1\]

OpenStudy (amistre64):

if the sequence does not limit to zero, the series is not convergent

OpenStudy (anonymous):

lol need not perform that much algebraic manipulation to see it diverges, clearly the terms are roughly one, so the sum grows linearly, and clearly diverges

OpenStudy (amistre64):

lol, but i like to practice archaic devices

OpenStudy (amistre64):

1-2/3 + 1 - 2/4 + 1 - 2/5 + 1 - 2/6 + .... your always adding 1 and subtracting less than 1 ... subtracting even smaller amounts less than 1 its not going to converge

OpenStudy (amistre64):

ahh, a sequence is not a series

OpenStudy (anonymous):

I've attached the problem, so I'm with you guys, it is divergent, but did I miss anything?

OpenStudy (amistre64):

the sequence converges to 1

OpenStudy (amistre64):

if the limit of the sequence is finite, then the sequence itself converges if the limit of the sequence is not 0, then the series (sum of the sequence) diverges

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ n }{ n+2 }=1\] so diverges

OpenStudy (anonymous):

I forget, what is the distinction between a sequence and a series?

OpenStudy (amistre64):

a sequence is a list of numbers a series as the sum of that list

OpenStudy (anonymous):

if sum is conv. then limit must be 0.... if limit is different than 0 then it is div..

OpenStudy (amistre64):

1,2,3,4,5,... is a sequence of numbers; a line up 1 + 2 + 3 + 4 + 5 + .... is a series (a sum of the numbers in the list)

OpenStudy (anonymous):

Oh, okay. Thank you very much for the help. Crystal clear.

OpenStudy (amistre64):

youre welcome